# Scanning gate microscopy probing of anisotropic electron flow in a two dimensional electron gas at the (110) $\mathrm{LaAlO}_{3}/\mathrm{SrTiO}_{3}$ interface

### v1

@@ I Introduction
The most commonly studied 2DEG in oxide heterostructures is the one created at the LAO/STO interface grown in [001] direction. The band structure of this interface, in the low-energy regime, is dominated by a parabolic and isotropic titanium $d_{xy}$ band separated by $50-300$ meV from highly anisotropic $d_{yz/xz}$ bands [12, 13]. Recently, a high mobility 2DEG has been realized at the LAO/STO (110) interface. Interestingly, it was found that the electron transport properties are highly anisotropic when comparing the two in-plane crystallographic directions [14]. A similar effect was demonstrated for the case of SrTiO3 (110) without a LaAlO3 layer but covered by a two-dimensional tetrahedrally coordinated titania overlayer [15]. In that case, ARPES measurements show two elliptical Fermi surfaces rotated with respect to each other by 90 degrees.
@@ II.1 Tight-binding model
$\begin{split}E^{xy}_{\mathbf{k},n}&=-4t_{1}-2t_{2}+2t_{2}\cos{k_{Z}a_{z}}\\ &-2\left[2t^{2}_{1}+2t^{2}_{1}\cos{(k_{M}a_{M})}\right]^{1/2}\cos{\left(\frac{\pi n}{N+1}\right)},\\ E^{yz/xz}_{\mathbf{k},n}&=-4t_{1}-2t_{2}+2t_{1}\cos{k_{Z}a_{Z}}\\ &-2\left[t^{2}_{1}+t^{2}_{2}+2t_{1}t_{2}\cos{(k_{M}a_{M})}\right]^{1/2}\cos{\left(\frac{\pi n}{N+1}\right)},\\ \end{split}$ (2)
and $xy$, $yz$, and $xz$ correspond to the three $t_{2g}$ orbitals at each lattice site, $N$ to the number of layers, and $n$ indexes the quantized energy levels that arise from confinement in the $[110]$ direction. The DFT calculations show that the inter-orbital terms are either zero or negligible, and the hopping parameters have the values $t_{1}=-0.277$ eV, $t_{2}=-0.031$ eV [15] leading to a degenerate $xz$ and $yz$ band. The two components of the momentum vector $\mathbf{k}=(k_{Z},k_{M})$ correspond to the two directions in momentum space defined by in-plane reciprocal lattice vectors, while $a_{M}=a$, $a_{Z}=\sqrt{2}a$, where $a=0.39$ nm is the distance between two Ti nearest neighboring lattice sites (cf. Fig. 1).
@@
$\begin{split}\epsilon_{\mathbf{k}l}&=2t^{l}_{Z}\big{(}\cos{(k_{Z}a_{Z})}-1\big{)}\\ &+2t^{l}_{M}\big{(}\cos{(k_{M}a_{M})}-1\big{)}+\delta_{l}.\\ \end{split}$ (6)
@@
By applying the back-gate voltage, one can tune the actual Fermi energy, which allows to access both the single- or two-band regime in the considered system as proposed in Ref. [32]. It should be noted that both the anisotropic effective hopping parameters and the band offset depend significantly on the thickness of the 2DEG only in the range $N\approx 1-10$ as shown in Fig. 2 (b) and (c). Moreover, the first $yz/xz$ excited state of the [110] quantization is always above the $xy$-ground state, meaning that $\Delta^{\prime}>0$, where
@@
Figure 2: (a) Low-energy bands of the (110) LAO/STO interface plotted along the high symmetry directions. The solid lines correspond to the effective lattice model given by Eq. (3) while the dots correspond to the original dispersion relations determined by Eq. (2) for $n=1$; (b) The effective hopping parameters as a function of the number of Ti layers in the [110] direction; (c) The $N$-dependence of the band offset between the $xy$ and $yz/xz$ bands for $n=1$ [$\Delta$ defined by Eq. (7)] as well as the energy difference between the first $yz/xz$-excited state and the $xy$-ground state [$\Delta^{\prime}$ defined by Eq. (8)] of the [110] quantization.
It should be noted that the effective model given by Eq. (3) is defined on a rectangular lattice with lattice constants: $a_{M}=a$, $a_{Z}=\sqrt{2}a$, where $a=0.39$ nm. This imposes a further limitation when it comes to simulations of devices of realistic size. To overcome this limitation, we apply a scaling procedure to the effective Hamiltonian, which allows us to reduce the number of lattice sites without reducing the size of the system [34]. Namely, we rescale our lattice constants by taking: $\tilde{a}^{\prime}_{M}=sa_{M}$, $\tilde{a}_{Z}=sa_{Z}$, where $s$ is the scaling parameter. As a consequence, the hopping parameters also have to be modified accordingly: $\tilde{t}^{l}_{Z}=t_{Z}^{l}\;(a_{Z}/\tilde{a}_{Z})^{2}$, $\tilde{t}^{l}_{M}=t_{Z}^{l}\;(a_{M}/\tilde{a}_{M})^{2}$.
@@
The potential $V_{SGM}(x,y)$ induced by the SGM tip at the level of the 2DEG is typically of Gaussian or Laplacian form [36, 37, 38, 39]. We adopt the latter potential distribution following Ref. [40]
@@ III.1 Anisotropic band structure
In Fig. 3 we depict three Fermi surfaces calculated for $E_{F}=2,5$ and $8$ meV, with colors representing the Fermi wavelength $\lambda=2\pi/|\mathbf{k}|$. The Fermi contours are distinctly non-circular, exhibiting a significant variation in the Fermi wavelengths along the contour. In the following, we will consider the case of $E_{F}=5$ meV.
@@ III.2 SGM conductance maps
Another surprising feature is the visible shift of the single and double lobes shown in Figs. 6(b) and (c) from the diagonal line extending from the QPC slit. This can be better observed in Fig. 7(a) obtained without the SGM tip potential. There, in gray, we denote the QPC gate potential, and in green, the probability density distribution for the QPC tuned to the first conductance step. The electrons escaping the constriction do not follow the $x=-y$ line pointing from the constriction; rather, they are bent upwards, which translates into the skewed lobes observed in Figs. 6(b) and (c). This phenomenon can be understood by inspecting the Fermi velocities. In Fig. 7(b) we show the Fermi contour along with the Fermi velocities calculated numerically from the energies given by the tight-binding model Eq. (9) with $\mathbf{v}=[\partial E/\partial k_{x},\partial E/\partial k_{y}]/\hbar$. We clearly observe that the $x$ component of the velocities is dominant as $m_{x}<m_{y}$, as also evidenced by the effective mass model approximating the velocities as $\mathbf{v}=\hbar[k_{x}/m_{x},k_{y}/m_{y}]$. In the QPC constriction, the electron wavefunction takes the form of a standing wave between the QPC electrodes and the expectation value of the momentum in the $x=y$ direction is $\langle k_{x}+k_{y}\rangle=0$. Upon escaping the constriction and entering the open region, the point $k_{x}+k_{y}=0$ on the Fermi surface corresponds to a higher velocity in the $x$ direction compared to the $y$ direction [see the orange dashed line in Fig. 7(b)]. As a result, the particle follows a trajectory oriented towards the positive $y$ direction when $m_{x}<m_{y}$.
@@ III.4 Weak constriction in [110] direction
The situation is different for the diagonal orientation of the QPC shown in Fig. 10. First of all, since the electrons in both bands possess a similar wave vector in the $x=-y$ direction, we anticipate that both bands should contribute to the overall conductance with a comparable length of the conductance steps. In Fig. 10(a), the blue curve represents the contribution to the conductance from the $d_{yz/xz}$ bands, while the red curve indicates the conductance contribution from the $d_{xy}$ band. The band separation is reflected in the shift of the red conductance trace towards smaller $V_{g}$ potentials. The black curve depicts the overall conductance.
@@ VI Acknowledgments
This work is financed by the Horizon Europe EIC Pathfinder under the grant IQARO number 101115190 titled "Spin-orbitronic quantum bits in reconfigurable 2D-oxides" and partly supported by the program „Excellence initiative – research university” for the AGH University. We gratefully acknowledge the Polish high-performance computing infrastructure PLGrid (HPC Center: ACK Cyfronet AGH) for providing computer facilities and support within computational grant no. PLG/2024/017374.

### v2

@@ I Introduction
The most commonly studied 2DEG in oxide heterostructures is the one created at the LAO/STO interface grown in [001] direction. The band structure of this interface, in the low-energy regime, is dominated by a parabolic and isotropic titanium $d_{xy}$ band separated by 5030050-300 meV from highly anisotropic $d_{yz/xz}$ bands [12, 13]. Recently, a high mobility 2DEG has been realized at the LAO/STO (110) interface. Interestingly, it was found that the electron transport properties are highly anisotropic when comparing the two in-plane crystallographic directions [14]. A similar effect was demonstrated for the case of SrTiO3 (110) without a LaAlO3 layer but covered by a two-dimensional tetrahedrally coordinated titania overlayer [15]. In that case, ARPES measurements show two elliptical Fermi surfaces rotated with respect to each other by 90 degrees. It was also proposed that electron diffraction can be used to elucidate the anisotropic dispersion in two-dimensional electron systems [16].
@@ II.1 Tight-binding model
[…]
and $xy$, yzyz, and $xz$ correspond to the three $t_{2g}$ orbitals at each lattice site, $N$ to the number of layers, and $n$ indexes the quantized energy levels that arise from confinement in the $[110]$ direction. The DFT calculations show that the inter-orbital terms are either zero or negligible, and the hopping parameters have the values $t_{1}=-0.277$ eV, $t_{2}=-0.031$ eV [15] leading to a degenerate $xz$ and $yz$ band. The two components of the momentum vector $\mathbf{k}=(k_{Z},k_{M})$ correspond to the two directions in momentum space defined by in-plane reciprocal lattice vectors, while $a_{M}=a$, $a_{Z}=\sqrt{2}a$, where $a=0.39$ nm is the distance between two Ti nearest neighboring lattice sites (cf. Fig. 1).
@@
[…]
@@
By applying the back-gate voltage, one can tune the Fermi energy [33], which allows accessing both the single- or two-band regime in the considered system as proposed in Ref. [34]. It should be noted that both the anisotropic effective hopping parameters and the band offset depend significantly on the thickness of the 2DEG only in the range $N\approx 1-10$ as shown in Fig. 2 (b) and (c). Moreover, the first $yz/xz$ excited state of the [110] quantization is always above the $xy$-ground state, meaning that $\Delta^{\prime}>0$, where
@@
Figure 2: (a) Low-energy bands of the (110) LAO/STO interface plotted along the high symmetry directions. The solid lines correspond to the effective lattice model given by Eq. (3) while the dots correspond to the original dispersion relations determined by Eq. (2) for $n=1$ with $N=6$; (b) The effective hopping parameters as a function of the number of Ti layers in the [110] direction; (c) The NN-dependence of the band offset between the $xy$ and $yz/xz$ bands for $n=1$ $\Delta$ defined by Eq. (7)] as well as the energy difference between the first $yz/xz$-excited state and the $xy$-ground state $\Delta^{\prime}$ defined by Eq. (8)] of the [110] quantization.
For a relatively low thickness of 2DEG at the surface of (110) oriented STO, both ARPES measurements and DFT/tight-binding calculations reported in Ref. [15] show a band structure similar to that considered here in Fig. 2. Namely, in both cases, the bottom of the yz/xz band is shifted with respect to the xy band by about 305030-50 meV, and the values of the effective masses resulting from our model and those obtained theoretically and experimentally in Ref. [15] are in relatively good agreement, as shown in Table 1.
$m_{Z}^{xy}$ $m_{Z}^{yz/xz}$ $m_{M}^{xy}$ $m_{M}^{yz/xz}$
Theor. $m^{*}$ $8.1$ $0.9$ $1.0$ $5.0$
Theor. $m^{*}$ from Ref.[15] $8.2$ $0.6$ $0.6$ $4.7$
Expt. $m^{*}$ from Ref.[15] $9.7$ $0.67$ $0.74$ $6.1$
Table 1: Effective mass in the units of electron mass along the two high symmetry directions ($M$ and $Z$) for the $xy$ and $yz/xz$ bands as calculated with the use of the considered tight-binding effective model with $N=6$ (first row). For comparison we provide both the theoretical and experimental values from Ref. [15] (second and third row).
It should be noted that the effective model given by Eq. (3) is defined on a rectangular lattice with lattice constants: $a_{M}=\sqrt{2}a$, $a_{Z}=a$, where $a=0.39$ nm. This imposes a further limitation when it comes to simulations of devices of realistic size. To overcome this limitation, we apply a scaling procedure to the effective Hamiltonian, which allows us to reduce the number of lattice sites without reducing the size of the system [36]. Namely, we rescale our lattice constants by taking: $\tilde{a}^{\prime}_{M}=sa_{M}$, $\tilde{a}_{Z}=sa_{Z}$, where $s$ is the scaling parameter. As a consequence, the hopping parameters also have to be modified accordingly: $\tilde{t}^{l}_{Z}=t_{Z}^{l}\;(a_{Z}/\tilde{a}_{Z})^{2}$, $\tilde{t}^{l}_{M}=t_{Z}^{l}\;(a_{M}/\tilde{a}_{M})^{2}$.
@@
The potential $V_{SGM}(x,y)$ induced by the SGM tip at the level of the 2DEG is typically of Gaussian or Lorentzian form [38, 39, 40, 41]. We adopt the latter potential distribution following Ref. [42]
@@ III.1 Anisotropic band structure
In Fig. 3 we depict three Fermi surfaces calculated for $E_{F}=2,5$ and $8$ meV, with colors representing the Fermi wavelength $\lambda=2\pi/|\mathbf{k}|$. The Fermi contours are distinctly non-circular, exhibiting a significant variation in the Fermi wavelengths along the contour. In the following, we will consider the case of $E_{F}=5$ meV. The concrete choice of the Fermi energy has no qualitative effect on the results presented here, and it only affects the Fermi wavelength, so by that the obtained distance between the conductance fringes.
@@ III.2 SGM conductance maps
Another surprising feature is the visible shift of the single and double lobes shown in Figs. 6(b) and (c) from the diagonal line extending from the QPC slit. This can be better observed in Fig. 7(a) obtained without the SGM tip potential. There, in gray, we denote the QPC gate potential, and in green, the probability density distribution for the QPC tuned to the first conductance step. The electrons escaping the constriction do not follow the x=yx=-y line pointing from the constriction; rather, they are bent upwards, which translates into the skewed lobes observed in Figs. 6(b) and (c). This phenomenon can be understood by inspecting the Fermi velocities. In Fig. 7(b) we show the Fermi contour along with the Fermi velocities calculated numerically from the energies given by the tight-binding model Eq. (9) with $\mathbf{v}=[\partial E/\partial k_{x},\partial E/\partial k_{y}]/\hbar$. We clearly observe that the $x$ component of the velocities is dominant as $m_{x}<m_{y}$, as also evidenced by the effective mass model approximating the velocities as $\mathbf{v}=\hbar[k_{x}/m_{x},k_{y}/m_{y}]$. In the QPC constriction, the electron wavefunction takes the form of a standing wave between the QPC electrodes and the expectation value of the momentum in the $x=y$ direction is $\langle k_{x}+k_{y}\rangle=0$. Upon escaping the constriction and entering the open region, the point $k_{x}+k_{y}=0$ on the Fermi surface corresponds to a higher velocity in the $x$ direction compared to the $y$ direction [see the orange dashed line in Fig. 7(b)]. As a result, the particle follows a trajectory oriented towards the positive $y$ direction when $m_{x}<m_{y}$.
@@ III.4 Weak constriction in [110] direction
The situation is different for the diagonal orientation of the QPC shown in Fig. 10. First of all, since the electrons in both bands possess a similar wave vector in the x=yx=-y direction, we anticipate that both bands should contribute to the overall conductance with a comparable length of the conductance steps. In Fig. 10(a), the blue curve represents the contribution to the conductance from the $d_{yz/xz}$ bands, while the red curve indicates the conductance contribution from the $d_{xy}$ band. The band separation is reflected in the shift of the red conductance trace towards smaller $V_{g}$ potentials. The black curve depicts the overall conductance.
@@ VI Acknowledgments
This work is financed by the Horizon Europe EIC Pathfinder under the grant IQARO number 101115190 titled "Spin-orbitronic quantum bits in reconfigurable 2D-oxides" and partly supported by the program „Excellence initiative – research university” for the AGH University. We gratefully acknowledge the Polish high-performance computing infrastructure PLGrid (HPC Center: ACK Cyfronet AGH) for providing computer facilities and support within computational grant no. PLG/2025/018486.